# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v2

@@ 7.1 LV-Model (BFF) Histograms
Figure 20: Histograms of the Lotka–Volterra parameters $(\alpha,\beta,\delta,\gamma)$.
Figure 21: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
@@ 7.2 LV-model (QDF) Histograms
Figure 23: Histograms for predator-prey data of $(\alpha,\beta,\delta,\gamma)$ Parameters
Figure 24: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
[…]
7.3 GLM-Model-Lung Injury (BFF) Histograms
Figure 26: Histograms for GLM data of r, k, and p Parameters
7.4 GLM-Model-Lung Injury (QDF) Histograms
Figure 27: Histograms r,p, and k Parameters
7.5 GLM-Model-Mpox (BFF) Histograms
Figure 28: Histograms for GLM data of r, k, and p Parameters
7.6 GLM-Model-Mpox (QDF) Histograms
Figure 29: Histograms for GLM data of r, k, and p Parameters
7.7 SEIUR-Model (BFF) Histograms
Figure 30: Histograms ($\beta_{0}$, $\beta_{1}$, $q$, rhorho, kappakappa, and $sigma_{1}$ ) Parameters
7.8 SEIUR-Model (QDF) Histograms
Figure 31: Histograms for ($\beta_{0}$, $\beta_{1}$, $q$, $\rho$, $\kappa$, and $\gamma_{1}$ ) Parameters

### v3

@@ 7.1 LV-model (BFF) histograms
Figure 20: Histograms of the LV parameters $(\alpha,\beta,\delta,\gamma)$.
Figure 21: Histograms for predator data only with $(\alpha,\beta,\delta,\gamma)$ parameters.
@@ 7.2 LV-model (QDF) Histograms
Figure 23: Histograms for predator prey data with the $(\alpha,\beta,\delta,\gamma)$ parameters.
Figure 24: Histograms for predator data only with the $(\alpha,\beta,\delta,\gamma)$ Parameters.
Figure 25: Histograms for prey data only with the $(\alpha,\beta,\delta,\gamma)$ parameters.
7.3 GLM model lung injury (BFF) histograms
Figure 26: Histograms for GLM data with the $r$, $k$, and $p$ parameters.
7.4 GLM model lung injury (QDF) histograms
Figure 27: Histograms $r$, $p$, and $k$ parameters.
7.5 GLM model mpox (BFF) histograms
Figure 28: Histograms for GLM data with the $r$, $k$, and $p$ parameters.
7.6 GLM model mpox (QDF) histograms
Figure 29: Histograms for GLM data with the $r$, $k$, and $p$ parameters.
7.7 SEIUR model (BFF) histograms
Figure 30: Histograms $\beta_{0}$, $\beta_{1}$, $q$, rhorho, kappakappa, and $sigma_{1}$ parameters.
7.8 SEIUR model (QDF) histograms
Figure 31: Histograms for $\beta_{0}$, $\beta_{1}$, $q$, $\rho$, $\kappa$, and $\gamma_{1}$ parameters.